OpenStudy (wilder.monday):

PLEASE HELP 1. 3^5x = 42

OpenStudy (campbell_st):

so is the question \[3^{5x} = 42\]

OpenStudy (wilder.monday):

yes :)

OpenStudy (campbell_st):

ok... so the easiest thing to do is

OpenStudy (campbell_st):

take the log of both sides.... does that make sense..?

OpenStudy (wilder.monday):

that what my teacher tried to teach me but i don't understand how :/

OpenStudy (campbell_st):

ok... so the log law you need to know about is the one that deals with powers base \[\log(a^b) = b \times \log(a)\] so the log of a number involving a power can be written as the power times the log of the

OpenStudy (campbell_st):

so you would have \[\log(3^{5x}) = \log(42)\] any ideas on how to rewrite the left hand side..?

OpenStudy (wilder.monday):

5x log 3 ?

OpenStudy (campbell_st):

great so you have \[5x \times \log(3) = \log(42 log(3) and log(42) are just numbers...dividing both sides by log(3) and that will give you a value for 5x, then you can some for x. you make want to check to question to see if you need a decimal approximation or an exact value

OpenStudy (campbell_st):

oops should read \[5x \times \log(3) = \log(42)\] and it doesn't matter which type of logs you use... base e or base 10 you'll get the same answer

OpenStudy (campbell_st):

I'd expect the value of x is between .6 and .8.... probably around 0.7

OpenStudy (wilder.monday):

so i divide in my calculator log 3/log 42

OpenStudy (wilder.monday):

oh no wait its the other way around

OpenStudy (campbell_st):

no other way \[5x = \frac{\log(42)}{\log(3)}\] then divide the answer by 5

OpenStudy (wilder.monday):

and then i divided that by 5 and the answer i got was 0.6804...

OpenStudy (campbell_st):

makes sense to me yo check you can use \[3^{5 \times 0.6804}\ and see if you get 42

OpenStudy (wilder.monday):

yes its right. thank you for your help. mind helping with a few more?

OpenStudy (campbell_st):

well post them as new questions and if I'm still here I'll look at them. There are lots of people her who provide great help